To test whether whether or not the product of two vendors had a larger mean value than the other, a sample of 12 pieces was taken from each. The T test value was calculated as 1.85 ( with equal variance assumed). Using a level of significance of 0.01, one tail test, we can conclude that

a. There was no significant difference in the variation of the vendors

b. The T test value is less than the t critical

c. The null hypothesis is that the means are the same was rejected.
Answer #1

Degrees of freedom = 12 + 12 - 2 = 22

For 22 degrees of freedom and 1.85 test statistic with one tailed test:

p - Value = P(t(22) > 1.85) = 0.0389

This is greater than 0.01 significance level. So,

The T test value is less than the t critical

Option B is correct.